# WOUAF: Weight Modulation for User Attribution and Fingerprinting in Text-to-Image Diffusion Models

### v2

@@ Abstract
The rapid advancement of generative models, facilitating the creation of hyper-realistic images from textual descriptions, has concurrently escalated critical societal concerns such as misinformation. Although providing some mitigation, traditional fingerprinting mechanisms fall short in attributing responsibility for the malicious use of synthetic images. This paper introduces a novel approach to model fingerprinting that assigns responsibility for the generated images, thereby serving as a potential countermeasure to model misuse. Our method modifies generative models based on each user’s unique digital fingerprint, imprinting a unique identifier onto the resultant content that can be traced back to the user. This approach, incorporating fine-tuning into Text-to-Image (T2I) tasks using the Stable Diffusion Model, demonstrates near-perfect attribution accuracy with a minimal impact on output quality. Through extensive evaluation, we show that our method outperforms baseline methods with an average improvement of 11% in handling image post-processes. Our method presents a promising and novel avenue for accountable model distribution and responsible use.
@@ 3 Methods
Fingerprinting Time ($\downarrow$)
Attribution Acc ($\uparrow$)
CLIP-score ($\uparrow$)
FID ($\downarrow$)
Attribution Acc ($\uparrow$)
CLIP-score ($\uparrow$)
FID ($\downarrow$)
Original SD [39]
0.73
24.48
0.50
19.67
DAG [22]
8.4 hr
0.70
0.73
26.54
0.71
0.49
23.13
Stable Signature [10]
$<$ 1 min
0.99
0.73
24.55
0.98
0.50
20.02
WOUAF-conv
$<$ 1 sec
0.99
0.73
24.43
0.98
0.51
20.46
WOUAF-all
$<$ 1 sec
0.99 0.73 24.42 0.99 0.51 19.91
@@ 3.1 Preliminaries
Our approach utilizes the Stable Diffusion (SD) model, which functions within the latent space framework of an autoencoder. SD comprises two main elements: Firstly, an autoencoder is pre-trained on an extensive dataset of images. Its encoder, $\mathcal{E}(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$, converts an image $x\sim p_{data}$ into a latent representation $z=\mathcal{E}(x)$. The decoder, $\mathcal{D}(\cdot):\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}^{d_{x}}$, then reconstructs the original image from this latent representation, resulting in $\hat{x}=\mathcal{D}(z)$. The secondary element is a diffusion model, based on the U-Net architecture [40], represented as $\epsilon_{\theta}$. This model is adept at generating latent representations and can be conditioned using pre-trained text embeddings
@@ 4.1 Experiment Settings
Experimental Setting. We implement the weight modulation following the design specified in the source code of StyleGAN2-ADA [19]. Our mapping network $M$ is designed with a series of fully connected layers, wherein all experiments are conducted using a two-layer configuration. The dimension of the mapping network $d_{M}$ is set to be equal to $4*d_{\phi}$ across all experimental setups. Training is performed over 50K iterations with a batch size of 32 and a learning rate of $10^{-4}$ using AdamW optimizer [28]. To train robust models against image post-processing transformations, differentiable post-processes are necessary. To this end, we incorporate the Kornia library [38]. For Stable Signature [10], we utilize the official code provided by the authors. We note that its post-processing transformations are replaced with our version for fair comparison.
@@ 4.2 Fingerprint Capacity
To investigate this trade-off, we conduct an analysis with varying fingerprint dimensions, specifically $d_{\phi}$ values of 16, 32, 64, and 128. Tab. 2 presents the user attribution accuracy for each $d_{\phi}$ value. As shown in Tab. 2, attribution accuracy tends to decrease monotonically as $d_{\phi}$ increases.
Importantly, both our variant models achieve a near-perfect attribution accuracy of 0.99 for $d_{\phi}$ values of 16, 32, and 64. However, for $d_{\phi}=128$, WOUAF-all variant outperforms the WOUAF-conv variant. For a balanced comparison with existing methods, we choose $d_{\phi}=32$, which notably can support a substantial user base exceeding 4 billion $\approx 2^{32}$.
@@ 4.4 Attribution Analysis for Diverse Image Sources
We adopted the experimental setup from [49], compiling a dataset with different image types: authentic images from the MS-COCO test set [27], non-fingerprinted images from Stable Diffusion [39], and synthesized images from ProGAN [16], StyleGAN [17], and StyleGAN2 [18], with each category containing 1,000 samples. Given our extensive user database of 1 million entries, we set a threshold at $32*0.95\approx 30$ bits, aligning with our 0.99 attribution accuracy as shown in Tab. 1.
@@ 4.6 Robust User Attribution against Image Post-processes
where, $T(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ denotes the post-processing function. In the optimization process, we employ an objective function akin to the one detailed in Eq. 4, with $L_{\phi}$ substituted by $L_{\text{robust}}$.
@@ 6 Conclusion
Acknowledgment
This work is partially supported by the National Science Foundation under Grant No. 2038666, No. 2101052, and a grant from Meta AI Learning Alliance. The HuggingFace demo of this work is funded by Intel Corporation. The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.
@@ Appendix A Additional Related Work
Recent advancements in vector quantization and diffusion modeling have significantly enhanced text-to-image (T2I) generation, enabling the creation of hyper-realistic images from textual prompts [31, 41, 37, 39]. These T2I models have been effectively utilized in various tasks such as generating images driven by subject, segmentation, and depth cues [4, 5, 35, 11, 25]. However, the substantial size of these models presents a challenge for broader user adoption. Research efforts are focusing on enhancing model efficiency through knowledge distillation, step distillation, architectural optimization, and refining text-to-image priors [26, 42, 29, 36]. Amidst these technological advancements, ensuring the responsible usage of these powerful tools is a critical area of focus, which is the aim of our proposed method.
@@ Appendix C Additional Experimental Results
C.1 Attribution Accuracy and Image Quality
@@
C.2 Evaluating Generalizability Across Datasets
@@
Attribution Acc. ($\uparrow$)
Clip-score ($\uparrow$)
FID ($\downarrow$)
Attribution Acc. ($\uparrow$)
Clip-score ($\uparrow$)
FID ($\downarrow$)
Ours-conv.
0.99
0.73
24.23
0.99
0.51
19.71
Ours-all
0.99
0.73
24.41
0.99
0.51
19.46
C.3 Attribution Accuracy Across Various Generation Hyperparameters
@@
Steps
Attribution Acc. ($\uparrow$)
CLIP-score ($\uparrow$)
FID ($\downarrow$)
Steps
Attribution Acc. ($\uparrow$)
CLIP-score ($\uparrow$)
FID ($\downarrow$)
Original SD [39]
20
0.73
24.48
50
0.73
23.33
WOUAF-conv
15
0.99
0.73
24.63
45
0.99
0.73
23.28
20
0.99
0.73
24.43
50
0.99
0.73
23.35
25
0.99
0.74
24.14
55
0.99
0.73
23.31
WOUAF-all
15
0.99
0.73
24.65
45
0.99
0.73
23.34
20
0.99
0.73
24.42
50
0.99
0.73
23.29
25
0.99
0.73
24.11
55
0.99
0.73
23.26
@@
Scheduler
Attribution Acc. ($\uparrow$)
CLIP-score ($\uparrow$)
FID ($\downarrow$)
Scheduler
Attribution Acc. ($\uparrow$)
CLIP-score ($\uparrow$)
FID ($\downarrow$)
WOUAF-conv
Euler
0.99
0.72
18.63
Euler
0.99
0.73
21.91
DDIM
0.99
0.72
18.35
DDIM
0.99
0.73
20.78
WOUAF-all
Euler
0.99
0.71
18.64
Euler
0.99
0.73
21.89
DDIM
0.99
0.71
18.31
DDIM
0.99
0.73
20.68
@@
C.4 Benefits of Finetuning only Decoder
@@
C.5 Robust User Attribution against Image Post-processes
Figure 12: Examples of images after applying each of the post-processing methods that we consider in the paper.
We conducted a thorough evaluation of quality metrics to assess the impact of our robust user attribution training on various image post-processing methods. Examples of images post-processed using these methods are displayed in Fig. 12. As indicated in Tab. 6 and Tab. 7, our robust fine-tuning approach generally preserves image quality with only minimal perturbations. A representative example under a JPEG attack, generated by our robust model, is showcased in Fig. 13. Additionally, our method demonstrates adaptability under Combination attacks, which significantly challenge image fidelity. As illustrated in Fig. 12, these combined post-processing techniques necessitate a relatively stronger fingerprint compared to single post-processes, as further detailed in Fig. 14.
@@ C.5 Robust User Attribution against Image Post-processes
Model
Crop
Rotation
Blur
Brightness
Noise
Erasing
JPEG
Combi.
WOUAF-conv
24.02
24.05
23.80
24.14
23.96
24.16
24.42
26.80
WOUAF-all
24.35
23.92
24.18
24.54
24.24
24.48
24.41
26.85
@@
Model
Crop
Rotation
Blur
Brightness
Noise
Erasing
JPEG
Combi.
WOUAF-conv
0.716
0.717
0.716
0.716
0.712
0.717
0.714
0.702
WOUAF-all
0.719
0.717
0.718
0.718
0.710
0.718
0.716
0.704

### v3

@@ Abstract
The rapid advancement of generative models, facilitating the creation of hyper-realistic images from textual descriptions, has concurrently escalated critical societal concerns such as misinformation. Although providing some mitigation, traditional fingerprinting mechanisms fall short in attributing responsibility for the malicious use of synthetic images. This paper introduces a novel approach to model fingerprinting that assigns responsibility for the generated images, thereby serving as a potential countermeasure to model misuse. Our method modifies generative models based on each user’s unique digital fingerprint, imprinting a unique identifier onto the resultant content that can be traced back to the user. This approach, incorporating fine-tuning into Text-to-Image (T2I) tasks using the Stable Diffusion Model, demonstrates near-perfect attribution accuracy with a minimal impact on output quality. Through extensive evaluation, we show that our method outperforms baseline methods with an average improvement of 11% in handling image post-processes. Our method presents a promising and novel avenue for accountable model distribution and responsible use. Our code is available in https://github.com/kylemin/WOUAF.
@@ 3 Methods
Fingerprinting Time ($\downarrow$) Attribution Acc ($\uparrow$) CLIP-score ($\uparrow$) FID ($\downarrow$) Attribution Acc ($\uparrow$) CLIP-score ($\uparrow$) FID ($\downarrow$)
Original SD [40] - - 0.73 24.48 - 0.50 19.67
DAG [22] 8.4 hr 0.70 0.73 26.54 0.71 0.49 23.13
Stable Signature [10] $<$ 1 min 0.99 0.73 24.55 0.98 0.50 20.02
WOUAF-conv $<$ 1 sec 0.99 0.73 24.43 0.98 0.51 20.46
WOUAF-all $<$ 1 sec 0.99 0.73 24.42 0.99 0.51 19.91
@@ 3.1 Preliminaries
Our approach utilizes the Stable Diffusion (SD) model, which functions within the latent space framework of an autoencoder. SD comprises two main elements: Firstly, an autoencoder is pre-trained on an extensive dataset of images. Its encoder, $\mathcal{E}(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$, converts an image $x\sim p_{data}$ into a latent representation $z=\mathcal{E}(x)$. The decoder, $\mathcal{D}(\cdot):\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}^{d_{x}}$, then reconstructs the original image from this latent representation, resulting in $\hat{x}=\mathcal{D}(z)$. The secondary element is a diffusion model, based on the U-Net architecture [41], represented as $\epsilon_{\theta}$. This model is adept at generating latent representations and can be conditioned using pre-trained text embeddings.
@@ 4.1 Experiment Settings
Experimental Setting. We implement the weight modulation following the design specified in the source code of StyleGAN2-ADA [19]. Our mapping network $M$ is designed with a series of fully connected layers, wherein all experiments are conducted using a two-layer configuration. To train robust models against image post-processing transformations, differentiable post-processes are necessary. To this end, we incorporate the Kornia library [39]. For Stable Signature [10], we utilize the official code provided by the authors. We note that its post-processing transformations are replaced with our version for fair comparison. Appendix includes details on mapping network dimensions, training parameters, and optimizer.
@@ 4.2 Fingerprint Capacity
To investigate this trade-off, we conduct an analysis with varying fingerprint dimensions, specifically $d_{\phi}$ values of 16, 32, 64, and 128. Tab. 2 presents the user attribution accuracy for each $d_{\phi}$ value. As shown in Tab. 2, attribution accuracy tends to decrease monotonically as $d_{\phi}$ increases. Importantly, both our variant models achieve a near-perfect attribution accuracy of 0.99 for $d_{\phi}$ values of 16, 32, and 64. However, for $d_{\phi}=128$, WOUAF-all variant outperforms the WOUAF-conv variant. For a balanced comparison with existing methods, we choose $d_{\phi}=32$, which notably can support a substantial user base exceeding 4 billion $\approx 2^{32}$.
@@ 4.4 Attribution Analysis for Diverse Image Sources
We adopt the experimental setup from [50], compiling a dataset with different image types: authentic images from the MS-COCO test set [27], non-fingerprinted images from Stable Diffusion [40], and synthesized images from ProGAN [16], StyleGAN [17], and StyleGAN2 [18], with each category containing 1,000 samples. Given our extensive user database of 1 million entries, we set a threshold at $32*0.95\approx 30$ bits, aligning with our 0.99 attribution accuracy as shown in Tab. 1.
@@ 4.6 Robust User Attribution against Image Post-processes
where $T(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ denotes the post-processing function. In the optimization process, we employ an objective function akin to the one detailed in Eq. 4, with $L_{\phi}$ substituted by $L_{\text{robust}}$.
@@ 6 Conclusion
7 Acknowledgment
This work is partially supported by the National Science Foundation under Grant No. 2038666, No. 2101052, and a grant from Meta AI Learning Alliance. The HuggingFace demo of this work is funded by Intel Corporation. The authors also acknowledge Research Computing at Arizona State University for providing HPC resources and support for this work. The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.
@@ Appendix A Additional Related Work
Recent advancements in vector quantization and diffusion modeling have significantly enhanced text-to-image (T2I) generation, enabling the creation of hyper-realistic images from textual prompts [31, 42, 38, 40]. These T2I models have been effectively utilized in various tasks such as generating images driven by subject, segmentation, and depth cues [4, 5, 35, 37, 11, 25]. However, the substantial size of these models presents a challenge for broader user adoption. Research efforts are focusing on enhancing model efficiency through knowledge distillation, step distillation, architectural optimization, and refining text-to-image priors [26, 43, 29, 36]. Amidst these technological advancements, ensuring the responsible usage of these powerful tools is a critical area of focus, which is the aim of our proposed method.
@@ Appendix C Additional Experimental Results
C.1 Additional Training Details
The dimension of the mapping network $d_{M}$ is set to be equal to $4*d_{\phi}$ across all experimental setups. Training is performed over 50K iterations with a batch size of 32 and a learning rate of $10^{-4}$ using AdamW optimizer [28].
C.2 Attribution Accuracy and Image Quality
@@
C.3 Evaluating Generalizability Across Datasets
@@
Attribution Acc. ($\uparrow$) Clip-score ($\uparrow$) FID ($\downarrow$) Attribution Acc. ($\uparrow$) Clip-score ($\uparrow$) FID ($\downarrow$)
Ours-conv. 0.99 0.73 24.23 0.99 0.51 19.71
Ours-all 0.99 0.73 24.41 0.99 0.51 19.46
C.4 Attribution Accuracy Across Various Generation Hyperparameters
@@
Steps Attribution Acc. ($\uparrow$) CLIP-score ($\uparrow$) FID ($\downarrow$) Steps Attribution Acc. ($\uparrow$) CLIP-score ($\uparrow$) FID ($\downarrow$)
Original SD [40] 20 - 0.73 24.48 50 - 0.73 23.33
WOUAF-conv 15 0.99 0.73 24.63 45 0.99 0.73 23.28
20 0.99 0.73 24.43 50 0.99 0.73 23.35
25 0.99 0.74 24.14 55 0.99 0.73 23.31
WOUAF-all 15 0.99 0.73 24.65 45 0.99 0.73 23.34
20 0.99 0.73 24.42 50 0.99 0.73 23.29
25 0.99 0.73 24.11 55 0.99 0.73 23.26
@@
Scheduler Attribution Acc. ($\uparrow$) CLIP-score ($\uparrow$) FID ($\downarrow$) Scheduler Attribution Acc. ($\uparrow$) CLIP-score ($\uparrow$) FID ($\downarrow$)
WOUAF-conv Euler 0.99 0.72 18.63 Euler 0.99 0.73 21.91
DDIM 0.99 0.72 18.35 DDIM 0.99 0.73 20.78
WOUAF-all Euler 0.99 0.71 18.64 Euler 0.99 0.73 21.89
DDIM 0.99 0.71 18.31 DDIM 0.99 0.73 20.68
@@
C.5 Benefits of Finetuning only Decoder
@@
C.6 Robust User Attribution against Image Post-processes
Figure 12: Qualitative results following the application of post-processing: These images demonstrate that intensive post-processing significantly compromises the perceptual quality of images, thereby deterring both malicious and benign users from employing robust post-processing techniques.
We conducted a thorough evaluation of quality metrics to assess the impact of our robust user attribution training on various image post-processing methods. Examples of images post-processed using these methods are displayed in Fig. 12. As indicated in Tab. 7 and Tab. 8, our robust fine-tuning approach generally preserves image quality with only minimal perturbations. A representative example under a JPEG attack, generated by our robust model, is showcased in Fig. 13. Additionally, our method demonstrates adaptability under Combination attacks, which significantly challenge image fidelity. As illustrated in Fig. 12, these combined post-processing techniques necessitate a relatively stronger fingerprint compared to single post-processes, as further detailed in Fig. 14. Moreover, it is observed that images subjected to extensive post-processing lose perceptual value, impacting both malicious and naive users alike.
C.7 Evaluating Robustness Beyond Training Configurations
Table 6: Assessment of generalizability of robust WOUAF. We measure the attribution accuracy for different models.
Crop Rotate Blur Bright Noise Erase
Params 30% $\pm$ 40 9 $\pm$ 0.4 0.3 30%
Stable Signature [10] 0.997 0.809 0.873 0.996 0.737 0.986
WOUAF-all 0.946 0.923 0.657 0.975 0.970 0.991
In Sec. 4.6 and Sec. C.6, we explore the resilience of our methodology against a variety of image post-processing techniques. Our experiments are designed with predefined post-process strengths sufficient to deter both malicious and benign uses of strong post-process modifications (refer to Fig. 12). Nonetheless, it is imperative to evaluate the resilience of our approach and the established baselines under conditions involving more intense attacks than those encountered during the training phase. In Tab. 6, we execute a series of tests to measure the attribution accuracy under more severe perturbations than those used in training and compare these results against those of baseline methodologies.
@@ C.7 Evaluating Robustness Beyond Training Configurations
Model Crop Rotation Blur Brightness Noise Erasing JPEG Combi.
WOUAF-conv 24.02 24.05 23.80 24.14 23.96 24.16 24.42 26.80
WOUAF-all 24.35 23.92 24.18 24.54 24.24 24.48 24.41 26.85
@@
Model Crop Rotation Blur Brightness Noise Erasing JPEG Combi.
WOUAF-conv 0.716 0.717 0.716 0.716 0.712 0.717 0.714 0.702
WOUAF-all 0.719 0.717 0.718 0.718 0.710 0.718 0.716 0.704
@@
Figure 15: Qualitative results of the original and fingerprinted Stable Diffusion models (WOUAF-conv) on MS-COCO. Pixel-wise differences are multiplied by a factor of 5 for a better view.